I need help with 1/3(7x+5)=3x-5.
Answer by rfer(16322)

7/3x+5/3=3x-5
7/3x-3x=-5-5/3
7/3x-9/3x=-15/3-5/3
-2/3x=-20/3
x=(-20/3)/(-2/3)
x=-20/3*-3/2
x=60/6
x=10